How many liters of each .5% saline solution and 2% saline to produce 10 liters of .9% saline?
I tried this using decimals but it came out to a negative number. I have the answer in the book but it is in fractions. Thank you!

Answer by lwsshak3(11628)

How many liters of each .5% saline solution and 2% saline to produce 10 liters of .9% saline?
***
let x=amt of .5% solution
10-x=amt of 2% solution
..
.5%x+2%(10-x)=.9%*10
.005x+.02(10-x)=.009*10
.005x+.20-.02x=.09
.015x=110
x=110/.015=22/3
10-x=30/3-22/3=8/3
amt of .5% solution=22/3 liters
amt of 2% solution=8/3 liters